# A coupled finite element-virtual element method for thermomechanical analysis of electronic packaging structures

### v1

@@ 5.3 Simplified sintered silver model
The third numerical example investigates a sintered silver interconnect structure commonly employed in high-power electronic packaging applications. As illustrated in Fig. 10, the model adopts a typical ‘sandwich’ configuration consisting of three distinct layers: a SiC chip (top), sintered silver interconnect layer (middle), and copper substrate (bottom). This multilayer structure is representative of modern power electronic modules where effective thermal management and mechanical reliability are critical design considerations [yu2021thermal]. The geometric dimensions are specified as follows: the SiC chip has width $b=1.8$ mm and thickness $h_{1}=0.5$ mm; the sintered silver layer maintains the same width as the SiC chip with thickness $h_{2}=0.3$ mm; and the copper substrate has width $a=3$ mm and thickness $h_{3}=0.8$ mm. The material properties for each component are detailed in Table 3, which reveal significant differences in thermal and mechanical characteristics that contribute to complex thermomechanical stress development during thermal loading.
@@
The top surface is subjected to a temperature of $T_{1}=150$ C representing heat generation from the semiconductor device, while the bottom surface is maintained at ambient temperature $T_{2}=25$ C. All lateral surfaces are treated as thermally insulated. Mechanically, the right edge is fully constrained ($u_{x}=u_{y}=0$). To demonstrate the capabilities of the FE-VE coupling approach, the upper layers (SiC chip and sintered silver) are discretized using virtual elements, while the copper substrate employs conventional finite elements.
Figure 11: Hybrid mesh discretization strategy for FEM-VEM coupling analysis.
For the numerical computation, the model is discretized using Abaqus preprocessing capabilities combined with custom MATLAB code, where the virtual element domain (upper layers) employs polygonal mesh elements to leverage VEM’s geometric flexibility, while the finite element domain (copper substrate) utilizes conventional four-node quadrilateral elements for computational efficiency. A critical aspect of the implementation is ensuring nodal compatibility at the coupling interface, where nodes from both domains must be properly matched to maintain solution continuity, as illustrated in Fig. 11. The figure demonstrates the distinct meshing approaches for each domain and the interface treatment, where polygonal elements transition to structured quadrilateral elements while maintaining nodal correspondence. This interface treatment ensures that displacement and stress continuity conditions are satisfied across the domain boundary, which is essential for the accuracy and stability of the FE-VE coupling scheme.
@@
The model has geometric parameters $a=40$ mm and $b=20$ mm and consists of three distinct parts with material properties listed in Table 4. Temperature boundary conditions are applied at the top surface ($T_{1}=150$ C) and bottom surface ($T_{0}=25$ C), with all other surfaces thermally insulated. The bottom surface is subjected to a fully fixed mechanical constraint.
@@ 5.4 FC-BGA Model
The fourth numerical example examines the thermomechanical performance of a Flip-Chip Ball Grid Array (FC-BGA) packaging structure under thermal loading conditions. FC-BGA packaging technology offers advantages of high integration density and superior performance with excellent product reliability and stability, making it widely adopted in high-performance processors and advanced semiconductor devices. As illustrated in Fig. 18, the computational model consists of six distinct components: mold compound, silicon die, adhesive epoxy, BT substrate, solder balls, and FR-4 PCB. The geometric dimensions of each component are detailed in Table 5.
@@
In this numerical example, the regular geometrical regions consisting of the BT substrate and FR-4 PCB are discretized using the finite element domain, while the remaining components are assigned to the virtual element domain. The finite element domain employs four-node quadrilateral elements for discretization, while the virtual element domain utilizes regular polygonal elements at interface boundaries and arbitrary polygonal elements in non-interface regions, with non-conforming mesh techniques used to integrate the VEM domain meshes into a unified structure. To ensure nodal compatibility at the coupling interface, careful attention is given to maintaining nodal correspondence between domains. Fig. 19 illustrates the final mesh discretization scheme, demonstrating the hybrid meshing strategy where complex geometries of the upper package components (mold compound, silicon die, adhesive epoxy) and solder balls are effectively handled by the flexible VEM approach, while structured lower layers benefit from the computational efficiency of conventional FEM. The interface region clearly shows how nodal correspondence between VE and FE domains is maintained to ensure solution continuity across the coupling boundary, demonstrating the seamless integration of both numerical methods in this complex multi-material electronic packaging configuration.
@@
To further validate the computational accuracy of the FE-VE coupling algorithm, a refined finite element analysis with mesh densification (nDof = 162,444) was performed as the reference solution. The displacement and stress profiles were extracted along two critical lines $L_{1}$ and $L_{2}$ (given in Fig. 18), as presented in Figs. 20 and 21. Fig. 20 demonstrates excellent agreement between the coupling algorithm and the reference FEM solution across all three response quantities along interface $L_{1}$. Fig. 20(a) shows the horizontal displacement exhibiting a nearly linear variation from approximately -0.003 mm to 0.004 mm, indicating consistent thermal expansion behavior with virtually identical results across all mesh refinement levels. Fig. 20(b) presents the vertical displacement distribution with a characteristic bell-shaped profile peaking at approximately 0.0065 mm at the interface center, corresponding to maximum thermal expansion due to the temperature gradient across package layers. Fig. 20(c) illustrates the stress distribution with periodic variations, showing peak stresses of approximately 60 MPa at interface boundaries and local minima of 44-46 MPa in central regions.
@@
Fig. 21 presents the von Mises stress distribution along line $L_{2}$ (shown in Fig. 18), which passes through the solder ball array. The stress profile exhibits pronounced periodic variations corresponding to the solder ball locations, with peak stresses reaching approximately 130 MPa at the solder ball interfaces and decreasing to about 20 MPa in the substrate regions between balls. The comparison between FE-VE coupling (nDof = 42,756) and reference FEM solution (nDof = 162,444) shows reasonable agreement in capturing the periodic stress pattern, though some differences are observed at stress concentration locations due to the coarser discretization in the coupling method. Despite the lower mesh density, the FE-VE approach successfully captures the overall stress distribution characteristics and solder ball interface behavior.
@@
Fig. 22 presents the nodal stress distribution along cross-material line $L_{3}$ (the blue line in Fig. 18), which traverses multiple material interfaces. The FE-VE coupling algorithm shows close agreement with the reference finite element solution across different material regions. Both methods capture the sharp stress concentration at $y\approx 1.86$ mm, where the peak stress reaches approximately 260 MPa due to significant material property mismatches at the interface. The stress distribution ranges from approximately 50 MPa in outer regions to the 260 MPa peak, demonstrating that the coupling algorithm maintains solution continuity when transitioning between materials with different thermal and mechanical properties. These results confirm the capability of the FE-VE methodology for multi-material thermomechanical problems in electronic packaging applications.
@@ 5.5 IGBT Module
Material parameters are assigned to each component as listed in Table 9, which reveals significant variations in thermal and mechanical characteristics that contribute to complex thermomechanical interactions. The boundary conditions simulate realistic operating conditions: a heat flux of $q=1000$ mW/mm2 is applied to the upper surface of the IGBT chip to represent power dissipation during operation, the lower surface of the Cu baseplate is maintained at $T=25^{\circ}$C representing heat sink cooling, and all remaining surfaces are treated as thermally insulated with $\nabla T\cdot\mathbf{n}=0$. This configuration creates a realistic thermal gradient from the heat-generating chip through the multi-layer package structure to the cooled baseplate. The substantial differences in thermal expansion coefficients, ranging from 2.5 $\times$ 10-6 C-1 for IGBT chip to 22.4 $\times$ 10-6 C-1 for solder, generate significant thermal stresses that make this an excellent validation case for the FE-VE coupling methodology in handling complex multi-material electronic packaging with curved geometries and high thermal gradients.
@@
Here, the complex Al bonding wire components are assigned to the virtual element domain, where polygonal elements facilitate mesh generation for intricate curved geometries, while the Cu baseplate is also discretized using the virtual element domain to reduce element count and computational cost. The remaining components are assigned to the finite element domain, employing four-node quadrilateral elements with careful attention to ensuring shared nodes at coupling interfaces to maintain solution continuity. Fig. 25 illustrates the hybrid mesh discretization strategy, demonstrating how the approach effectively combines VEM’s geometric flexibility for complex wire profiles with FEM’s computational efficiency for regular substrate layers. The interface regions show where nodal compatibility between VEM and FEM domains is maintained, enabling seamless integration of both numerical methods in this challenging multi-material electronic packaging configuration.
@@
Fig. 26 presents the vertical displacement distribution contours for the Al bonding wire, showing close agreement between both methods. The displacement values range from 0 to approximately 0.0126 mm for the coupling method and 0 to 0.0127 mm for the FEM method, demonstrating consistent displacement patterns throughout the complex wire geometry due to thermal expansion effects.
@@
To provide detailed validation, nodal values were extracted along interface line $L_{1}$ (the blue line in Fig. 23) using refined finite element analysis (nDof = 419,250) as the reference solution, with results presented in Fig. 27. The coupling algorithm shows good consistency across all response quantities: horizontal displacement varies linearly from -2.0 $\times$ 10-3 mm to 1.0 $\times$ 10-3 mm, vertical displacement peaks at approximately 2.9 $\times$ 10-3 mm, and stress distribution exhibits sharp peaks reaching 200 MPa at critical interface locations.
Fig. 28 shows results along cross-material line $L_{2}$ (the red line in Fig. 23), where stress concentrations reach 60 MPa and vertical displacement varies from 1.8 $\times$ 10-4 mm to 2.9 $\times$ 10-4 mm. The coupling algorithm maintains consistency with the FEM reference solution across different mesh refinement levels, confirming the effectiveness of the proposed methodology for complex IGBT thermal stress analysis.
@@ 6 Conclusions
This FE-VE coupled framework advances computational mechanics by providing enhanced flexibility for analyzing electronic packaging technologies with complex geometric features. Future research directions include: (i) extending the approach to 3D thermomechanical problems, (ii) incorporating more sophisticated material models including plasticity and damage mechanics, (iii) investigating adaptive mesh refinement strategies for optimal domain partitioning, and (iv) developing fully coupled solutions for strongly bidirectional thermomechanical interactions. To facilitate reproducibility and further development, source codes and implementation examples are publicly available at https://github.com/yanpeng-gong/FeVeCoupled-ElectronicPackaging.

### v2

@@ 5.3 Simplified sintered silver model
The third numerical example investigates a sintered silver interconnect structure commonly employed in high-power electronic packaging applications. As illustrated in Fig. 10, the model adopts a typical ‘sandwich’ configuration consisting of three distinct layers: a SiC chip (top), sintered silver interconnect layer (middle), and copper substrate (bottom). This multilayer structure is representative of modern power electronic modules where effective thermal management and mechanical reliability are critical design considerations [yu2021thermal]. The geometric dimensions are specified as follows: the SiC chip has width $b=1.8$ mm and thickness $h_{1}=0.5$ $m$; the sintered silver layer maintains the same width as the SiC chip with thickness $h_{2}=0.3$ $m$; and the copper substrate has width $a=3$ mm and thickness $h_{3}=0.8$ $m$. The material properties for each component are detailed in Table 3, which reveal significant differences in thermal and mechanical characteristics that contribute to complex thermomechanical stress development during thermal loading.
@@
The top surface is subjected to a temperature of $T_{1}=150$ C representing heat generation from the semiconductor device, while the bottom surface is maintained at ambient temperature $T_{2}=25$ C. All lateral surfaces are treated as thermally insulated. Mechanically, the right edge is fully constrained ($u_{x}=u_{y}=0$) and other boundaries of the model are set as free boundaries.To demonstrate the capabilities of the FE-VE coupling approach, the upper layers (SiC chip and sintered silver) are discretized using virtual elements, while the copper substrate employs conventional finite elements.
Figure 11: Hybrid mesh discretization strategy for FE-VE coupling analysis
For the numerical computation, the model is discretized using Abaqus preprocessing capabilities combined with custom MATLAB code, where the virtual element domain (upper layers) employs polygonal elements to leverage VEM’s geometric flexibility, while the finite element domain (copper substrate) utilizes conventional four-node quadrilateral elements for computational efficiency. A critical aspect of the implementation is ensuring nodal compatibility at the coupling interface, where nodes from both domains must be properly matched to maintain solution continuity, as illustrated in Fig. 11. The figure demonstrates the distinct meshing approaches for each domain and the interface treatment, where polygonal elements transition to structured quadrilateral elements while maintaining nodal correspondence. This interface treatment ensures that displacement and stress continuity conditions are satisfied across the domain boundary, which is essential for the accuracy and stability of the FE-VE coupling scheme.
@@
The model has geometric parameters $a=40$ mm and $b=20$ mm and consists of three distinct parts with material properties listed in Table 4. Temperature boundary conditions are applied at the top surface ($T_{1}=150$ C) and bottom surface ($T_{0}=25$ C), with all other surfaces thermally insulated. The bottom surface is subjected to a fully fixed mechanical constraint and other boundaries of the model are set as free boundaries.
@@ 5.4 FC-BGA Model
The fourth numerical example examines the thermomechanical performance of a Flip-Chip Ball Grid Array (FC-BGA) packaging structure under thermal loading conditions. FC-BGA packaging technology offers advantages of high integration density and superior performance with excellent product reliability and stability, making it widely adopted in high-performance processors and advanced semiconductor devices. As illustrated in Fig. 18, the computational model consists of six distinct components: mold compound, silicon die, adhesive epoxy, BT substrate, solder balls, and FR-4 PCB. The geometric dimensions of each component are detailed in Table 5. The solder balls are uniformly distributed with a pitch of H=1.24 mm.
@@
In this numerical example, the regular geometrical regions consisting of the BT substrate and FR-4 PCB are discretized using the finite element domain, while the remaining components are assigned to the virtual element domain. The finite element domain employs four-node quadrilateral elements for discretization, while the virtual element domain utilizes regular polygonal elements at interface boundaries and arbitrary polygonal elements in non-interface regions, with non-matching mesh techniques used to integrate the VEM domain meshes into a unified structure. To ensure nodal compatibility at the coupling interface, careful attention is given to maintaining nodal correspondence between domains. Fig. 19 illustrates the final mesh discretization scheme, demonstrating the hybrid meshing strategy where complex geometries of the upper package components (mold compound, silicon die, adhesive epoxy) and solder balls are effectively handled by the flexible VEM approach, while structured lower layers benefit from the computational efficiency of conventional FEM. The interface region clearly shows how nodal correspondence between VE and FE domains is maintained to ensure solution continuity across the coupling boundary, demonstrating the seamless integration of both numerical methods in this complex multi-material electronic packaging configuration.
Figure 20: Temperature distribution comparison for FC-BGA.
Fig. 20 presents the temperature field distributions computed by the two methods. The FEM solution with mesh refinement (nDof = 162,444) is shown for comparison. Both methods demonstrate consistent temperature distributions ranging from 50 C to 500 C.
@@
To further validate the computational accuracy of the FE-VE coupling algorithm, a refined finite element analysis with mesh densification (nDof = 162,444) was performed as the reference solution. The displacement and stress profiles were extracted along two critical lines $L_{1}$ and $L_{2}$ (given in Fig. 18), as presented in Figs. 21 and 22. Fig. 21 demonstrates excellent agreement between the coupling algorithm and the reference FEM solution across all three response quantities along interface $L_{1}$. Fig. 21(a) shows the displacement $u_{x}$ exhibiting a nearly linear variation, indicating consistent thermal expansion behavior with virtually identical results across all mesh refinement levels. Fig. 21(b) presents a bell-shaped distribution of $u_{y}$, with the most significant thermal expansion effect at the center due to the temperature gradient. Fig. 21(c) illustrates the stress distribution with periodic variations, showing peak stresses occurring at the interface boundaries and lower stresses in the central region.
@@
(a) FE-VE coupling solution in nDof=42,756
(b) FEM reference solution in nDof=162,444
Figure 23: Temperature distribution comparison for FC-BGA.
Fig. 22 presents the von Mises stress distribution along line $L_{2}$ (shown in Fig. 18), which passes through the solder ball array. The stress profile exhibits pronounced periodic variations corresponding to the solder ball locations.Fig. 23 displays the von Mises stress distributions around the solder joints obtained using both methods. The comparison between FE-VE coupling (nDof = 42,756) and reference FEM solution (nDof = 162,444) shows reasonable agreement in capturing the periodic stress pattern, though some differences are observed at stress concentration locations due to the coarser discretization in the coupling method. Despite the lower mesh density, the FE-VE approach successfully captures the overall stress distribution characteristics and solder ball interface behavior.
@@
Fig. 24 presents the nodal stress distribution along cross-material line $L_{3}$ (the blue line in Fig. 18), which traverses multiple material interfaces. The FE-VE coupling algorithm shows close agreement with the reference finite element solution across different material regions. The stress distribution demonstrating that the coupling algorithm maintains solution continuity when transitioning between materials with different thermal and mechanical properties. These results confirm the capability of the FE-VE methodology for multi-material thermomechanical problems in electronic packaging applications.
@@ 5.5 IGBT Module
Material parameters are assigned to each component as listed in Table 9, which reveals significant variations in thermal and mechanical characteristics that contribute to complex thermomechanical interactions. The boundary conditions simulate realistic operating conditions: a heat flux of $q=1000$ mW/mm2 is applied to the upper surface of the IGBT chip to represent power dissipation during operation, the lower surface of the Cu baseplate is maintained at $T=25^{\circ}$C representing heat sink cooling, and all remaining surfaces are treated as thermally insulated with $\nabla T\cdot\mathbf{n}=0$. The bottom surface is subjected to a fully fixed mechanical constraint (ux = uy = 0) and other boundaries of the model are set as free boundaries. This configuration creates a realistic thermal gradient from the heat-generating chip through the multi-layer package structure to the cooled baseplate. The substantial differences in thermal expansion coefficients generate significant thermal stresses that make this an excellent validation case for the FE-VE coupling methodology in handling complex multi-material electronic packaging with curved geometries and high thermal gradients.
@@
Here, the complex Al bonding wire component is assigned to the virtual element domain, where polygonal elements facilitate mesh generation for intricate curved geometries, while the Cu baseplate is also discretized using the virtual element domain to reduce element count and computational cost. The remaining components are assigned to the finite element domain, employing four-node quadrilateral elements with careful attention to ensuring shared nodes at coupling interfaces to maintain solution continuity. Fig. 27 illustrates the hybrid mesh discretization strategy, demonstrating how the approach effectively combines VEM’s geometric flexibility for complex wire profiles with FEM’s computational efficiency for regular substrate layers. The interface regions show where nodal compatibility between VEM and FEM domains is maintained, enabling seamless integration of both numerical methods in this challenging multi-material electronic packaging configuration.
Figure 28: Temperature distribution comparison for IGBT module .
Fig. 28 displays the temperature field distributions under heat flux. Both methods show excellent agreement, with temperatures ranging from 25 C to 57.97 C.
@@
Fig. 29 presents the vertical displacement distribution contours for the Al bonding wire, showing close agreement between both methods. The displacement values range from 0 to approximately 0.01272 mm for the coupling method and 0 to 0.0127 mm for the FEM method, demonstrating consistent displacement patterns throughout the complex wire geometry due to thermal expansion effects.
@@
To provide detailed validation, nodal values were extracted along interface line $L_{1}$ (the blue line in Fig. 25) using refined finite element analysis (nDof = 419,250) as the reference solution, with results presented in Fig. 30. The coupling algorithm shows good consistency across all response quantities.
Fig. 31 shows results along cross-material line $L_{2}$ (the red line in Fig. 25), where stress concentrations reach 60 MPa and vertical displacement varies from 1.8 $\times$ 10-3 mm to 2.9 $\times$ 10-3 mm. The coupling algorithm maintains consistency with the FEM reference solution across different mesh refinement levels, confirming the effectiveness of the proposed methodology for complex IGBT thermal stress analysis.
@@ 6 Conclusions
This FE-VE coupled framework advances computational mechanics by providing enhanced flexibility for analyzing electronic packaging technologies with complex geometric features. Future research directions include: (i) extending the approach to 3D thermomechanical problems, (ii) incorporating more sophisticated material models including plasticity and damage mechanics, (iii) investigating adaptive mesh refinement strategies for optimal domain partitioning, and (iv) developing fully coupled solutions for strongly bidirectional thermomechanical interaction problems. To facilitate reproducibility and further development, source codes and implementation examples are publicly available at https://github.com/yanpeng-gong/FeVeCoupled-ElectronicPackaging.